# Quantum geometry and low-frequency optical conductivity of topological nodal planes

### v2

@@ V Optical conductivity of a tight-binding model
In order to study, whether the characteristic low-frequency behaviour is visible also in more complicated systems, we consider a generic tight-binding model exhibiting a single nodal plane. Specifically, we examine a two-band tight-binding model for a hexagonal space group, which was introduced in [42], see Appendix D. We modify its dispersion by adding the term $-\cos(k_{z})\mathbb{1}$ to increase the energy separation of the nodal plane and the Weyl points. This is crucial in order to study the signatures of the nodal plane alone, since, as shown in Sec. IV, Weyl points close to the Fermi level will likely dominate the optical response. We choose the chemical potential such that a broad region of the nodal plane lies close to the Fermi energy. The resulting band structure is shown in Fig. 3(a). It has a nodal plane at $k_{z}=\pi$, which carries a topological charge $\nu=2$, and two Weyl points at the K points, which compensate the nodal-plane charge. For the numerical evaluation of the optical conductivity, a phenomenological broadening $\Gamma$ is introduced, i.e., the delta function $\delta(\omega-\frac{\Delta E}{\hbar})$ is replaced with a Lorentzian function $\frac{\Gamma}{\pi((\omega-\frac{\Delta E}{\hbar})^{2}+\Gamma^{2})}$. The resulting interband linear optical conductivity $\sigma^{zz}$ is shown in Fig. 3(b) for a broadening of $\Gamma=0.02$. For low frequencies ($\omega\leq 0.5$) we show the exact analytical result obtained for the low-energy model using the parameters of the tight-binding model (for details see Appendix D), and we perform a fit of the form $x^{\alpha}$ to extract the power-law behaviour, which yields $\alpha=3.07\pm 0.01$, agreeing with our expectations of a (larger than) cubic behaviour. The inset in Fig. 3(b) shows the low-frequency behaviour using logarithmic scales on both axes. We find that the power-law behaviour agrees very well except for very small frequencies comparable to the size of the broadening $\Gamma=0.02$, where a deviation is expected. Further deviations of the numerical result of the tight-binding model and the analytical results for the low-energy model are caused by the non-uniform dispersion perpendicular to the nodal plane and the dispersion of the nodal plane itself in the tight-binding model. The low-energy model of the nodal plane is obtained by performing a series expansion around a single point of the nodal plane and can, therefore, not capture the exact quantitative behaviour. Qualitatively, however, we find that the low-frequency behaviour agrees very well. Therefore, for this simple tight-binding model with a dispersive topological nodal plane, we can explain the low-frequency features with the results we obtain for a low-energy nodal plane. In particular, we find a $\omega^{3}$ behaviour in the component perpendicular to the nodal plane. Let us note, however, that in order to understand, whether this signature can be seen in experiment, intraband contributions need to be considered as well. Depending on the size of the phenomenological broadening, the broadened Drude peak can overlap significantly with the low-frequency spectrum, making a reliable analysis of the behaviour challenging. Furthermore, we consider a model with a single nodal plane, which is particularly common in hexagonal systems [72]. In general, the symmetries enforcing nodal planes can lead to nodal plane duos or trios [11]. In these cases, in-plane contributions to the optical conductivity from the different nodal planes are expected to lead also to lower-order contributions in $\omega$.
VI Conclusion
@@
In this work we introduced a generic low-energy model for a topological nodal plane, which is enforced by $\tilde{\mathcal{C}}_{2}^{z}\mathcal{T}$ symmetry. We found three distinct cases depending on the choice of parameter $a$, which exhibit a flat topological nodal plane and, in case of $a<0$, two additional Weyl points in the band structure. We then studied the quantum geometry, i.e., the Berry curvature and quantum metric, of this model, which shows the connection of a topological nodal plane to a Weyl point in the case of $a=0$. While the optical response of Weyl semimetals is well understood, the characteristic fingerprints of nodal planes in optical responses has not been studied so far. Here, we tackled this question by analytically determining the interband linear optical conductivity of the low-energy model, where we find a characteristic $\omega^{3}$ leading-order behaviour for $a>0$. We expect this behaviour to show up generically in the conductivity component perpendicular to the nodal plane. To support this claim, we considered a two-band tight-binding model with a single dispersive topological nodal plane at the Fermi level. The linear optical response of this system agrees well with the predicted $\omega^{3}$ behaviour for low frequencies. This is a first step towards using optical responses to detect topological nodal planes in experiments. As a next step, calculating these responses for material candidates including intraband contributions is necessary to see whether the low-frequency behaviour is overlapped by the broadened Drude peak. Further, we found that this characteristic feature does not depend on a finite topological charge, but also shows up in a topologically trivial nodal plane low-energy model.
@@ Acknowledgements.
As described in the previous section, the non-symmorphic symmetry enforcing the nodal plane at $k_{z}=0$ restricts the terms in the Hamiltonian to be odd in $k_{z}$. One finds that the lowest order in the $zz$ component of the optical conductivity is $\sim\omega^{3}$. In the following, we show that this result is still valid if one includes an arbitrary (symmetry-allowed) dispersion into the nodal plane low-energy expansion of Eq. (1), i.e., $\mathcal{H}_{\text{disp}}(\mathbf{k})=\mathcal{H}(\mathbf{k})+d_{0}(\mathbf{k})\sigma_{0}$ with some function $d_{0}(\mathbf{k})$. As introduced in Eq. (8) we consider
@@
The extra term $d_{0}(\vec{k})\sigma_{0}$ neither influences $\vec{n}$ nor $\Delta E$. It only takes effect on $\Delta f$, which is equal to 1 in the main text where for a finite gap the Fermi energy lies between the two bands. Further, note that the evaluation of the $\delta(\omega-\frac{\Delta E}{\hbar})$ term results in the integral being reduced to a surface integration. Solving $\omega-\frac{\Delta E}{\hbar}=0$ for $k_{z}$, we find that due to the presence of the symmetry-enforced nodal plane $k_{z,0}=a(k_{x},k_{y})\omega+\mathcal{O}(\omega^{2})$ with some function $a(k_{x},k_{y})$, as long as the nodal plane is the only degeneracy of Eq. (1). Then similarly, we get the low frequency expansion $\partial_{z}\vec{n}\cdot\partial_{z}\vec{n}=b(k_{x},k_{y})\omega^{2}+\mathcal{O}(\omega^{3})$ by plugging in the previous $k_{z}$ expansion. With $d_{0}(\mathbf{k})=0$ (the case from the main text) and these considerations, Eq. (8) for $a=b=z$ becomes
@@
×$\times\int\text{d}k_{x}\text{d}k_{y}h(k_{x},k_{y})\Delta f(k_{x},k_{y},k_{z}=k_{z,0}(\omega))+\mathcal{O}(\omega^{4})\;.$ (16)
@@
×$\hskip 56.9055pt\times\left(63a^{2}+\frac{15a\left(f(\omega)-2a\varepsilon^{2}\right)^{2}}{\varepsilon^{2}f(\omega)}+\frac{35\left(f(\omega)-2a\varepsilon^{2}\right)^{4}}{36\varepsilon^{4}f(\omega)^{2}}\right)\;,$ (18)
@@
×$\hskip 56.9055pt\times 2t_{\mathrm{c}}\cos(k_{z})-\cos(k_{z})+\mu$
@@
×$\hskip 56.9055pt\times 2t_{0}\cos\left(\frac{k_{z}}{2}\right)$
@@
×$\hskip 56.9055pt\times 2t_{0}\cos\left(\frac{k_{z}}{2}\right)$
@@
×$\hskip 56.9055pt\times 4t_{\mathrm{c}}\sin\left(\frac{k_{x}}{2}\right)\sin(k_{z})\;,$
where $\sigma$ is the vector of Pauli matrices. We added the $-\cos(k_{z})$ term and the chemical potential $\mu$ in $d_{0}$ to move the Weyl points far below the Fermi energy such that their contribution to the optical conductivity is suppressed. For our calculation we use the parameters $t_{0}=0.4$ and $t_{\mathrm{c}}=0.1$ and we set the chemical potential to $\mu=-1.24$. The resulting band structure is shown in Fig. 3(a).
By expanding the Hamiltonian (19) around the high-symmetry point H $=(\frac{4\pi}{3},0,\pi)^{\mathrm{T}}$ one obtains the low-energy model
@@
[…]
which allows us to understand the dependence of the analytical solution on the parameters of the tight-binding model. Note, that for this we neglected further terms cubic in $k$, which are expected to only give small corrections in the prefactors. Furthermore, the energy offset caused by the dispersion and the chemical potential is neglected. This low-energy model corresponds to the $a=-1$ phase discussed for the model in Eq. (1), which means that two additional Weyl points are present. In order to remove their contribution to the optical response, we change the phase to correspond to the $a=+1$ case by introducing the parameter $b$ and setting $b=-1$. With this, the analytical behaviour shown in Fig. 3(b) is obtained.
@@
×$\times\left.\left(\frac{2}{3}\sqrt{\frac{2}{3}}\left(f(\omega)-\frac{20}{f(\omega)}-4\right)^{3/2}+\frac{\left(f(\omega)-\frac{20}{f(\omega)}-4\right)^{5/2}}{6\sqrt{6}}+\sqrt{6}\sqrt{f(\omega)-\frac{20}{f(\omega)}-4}\right)\right]\;,$ (23)
@@
which shows that this low-frequency behaviour does not depend on the topological charge but is a characteristic feature of nodal planes. In particular, every Berry-curvature component in this model vanishes, demonstrating again the important role of the quantum metric for the investigation of response properties of materials.

### v3

@@ VI Optical conductivity of a tight-binding model
For the numerical evaluation of the optical conductivity of this model, a phenomenological broadening $\Gamma$ is introduced, i.e., the delta function $\delta(\omega-\frac{\Delta E}{\hbar})$ is replaced with a Lorentzian function $\frac{\Gamma}{\pi((\omega-\frac{\Delta E}{\hbar})^{2}+\Gamma^{2})}$. The resulting interband linear optical conductivity $\sigma^{zz}$ is shown in Fig. 3(b) for a broadening of $\Gamma=0.02$. For low frequencies ($\omega\leq 0.5$) we show the exact analytical result obtained for the low-energy model using the parameters of the tight-binding model (for details see Appendix D), and we perform a fit of the form $x^{\alpha}$ to extract the power-law behaviour, which yields $\alpha=3.07\pm 0.01$, agreeing with our expectations of a (larger than) cubic behaviour. The inset in Fig. 3(b) shows the low-frequency behaviour using logarithmic scales on both axes. We find that the power-law behaviour agrees very well except for very small frequencies comparable to the size of the broadening $\Gamma=0.02$, where a deviation is expected. Further deviations of the numerical result of the tight-binding model and the analytical results for the low-energy model are caused by the non-uniform dispersion perpendicular to the nodal plane and the dispersion of the nodal plane itself in the tight-binding model. The low-energy model of the nodal plane is obtained by performing a series expansion around a single point of the nodal plane and can, therefore, not capture the exact quantitative behaviour. Qualitatively, however, we find that the low-frequency behaviour agrees very well. Therefore, for this simple tight-binding model with a dispersive topological nodal plane, we can explain the low-frequency features with the results we obtain for a low-energy nodal plane. In particular, we find a $\omega^{3}$ behaviour in the component perpendicular to the nodal plane.
VII Discussion and Material Candidates
@@
In this work we introduced a generic low-energy model for a topological nodal plane, which is enforced by $\tilde{\mathcal{C}}_{2}^{z}\mathcal{T}$ symmetry. We found three distinct cases depending on the choice of parameter $a$, which exhibit a flat topological nodal plane and, in case of $a<0$, two additional Weyl points in the band structure. We then studied the quantum geometry, i.e., the Berry curvature and quantum metric, of this model, which shows the connection of a topological nodal plane to a Weyl point in the case of $a=0$. While the optical response of Weyl semimetals is well understood, the characteristic fingerprints of nodal planes in optical responses has not been studied so far. Here, we tackled this question by analytically determining the interband linear optical conductivity of the low-energy model, where we find a characteristic $\omega^{3}$ leading-order behaviour for $a>0$. We expect this behaviour to show up generically in the conductivity component perpendicular to the nodal plane. To support this claim, we considered a two-band tight-binding model with a single dispersive topological nodal plane at the Fermi level. The linear optical response of this system agrees well with the predicted $\omega^{3}$ behaviour for low frequencies. Further, we found that this characteristic feature does not depend on a finite topological charge, but also shows up in a topologically trivial nodal plane low-energy model.
In order to understand, whether this signature can be seen in experiment, intraband contributions need to be considered as well. Depending on the size of the phenomenological broadening, the broadened Drude peak can overlap significantly with the low-frequency spectrum, making a reliable analysis of the behaviour challenging. Furthermore, we consider a model with a single nodal plane, which is particularly common in hexagonal systems [72]. In general, the symmetries enforcing nodal planes can lead to nodal plane duos or trios [11]. In these cases, in-plane contributions to the optical conductivity from the different nodal planes are expected to lead also to lower-order contributions in $\omega$.
Additional topological nodal features, such as Weyl points, can also influence the low-frequency optical signature if they are in the vicinity the Fermi level. Therefore, systems with a single trivial nodal plane are expected to be better suited since the bands forming the nodal plane are not required to exhibit additional crossings to compensate the topological charge. In Appendix F, we discuss the criteria for materials with which we expect the low-frequency signature to be visible most clearly and over a significant frequency range, we list all space groups with single nodal planes, and we provide suitable materials. In particular, we propose Sb3Sr5 [73] and XXMo3S3 ($X$ = Rb, K, Cs) [25, 74] to be promising materials for the measurement of the optical signature of nodal planes.
@@ Acknowledgements.
Data availability
The data that support the findings of this article are openly available [75].
@@
As described in the previous section, the non-symmorphic symmetry enforcing the nodal plane at $k_{z}=0$ restricts the terms in the Hamiltonian to be odd in $k_{z}$. One finds that the lowest order in the $zz$ component of the optical conductivity is $\sim\omega^{3}$. In the following, we show that this result is still valid if one includes an arbitrary (symmetry-allowed) dispersion into the nodal plane low-energy expansion of Eq. (1), i.e., $\mathcal{H}_{\text{disp}}(\mathbf{k})=\mathcal{H}(\mathbf{k})+d_{0}(\mathbf{k})\mathbb{1}$ with some function $d_{0}(\mathbf{k})$. As introduced in Eq. (8) we consider
@@
The extra term $d_{0}(\vec{k})\mathbb{1}$ neither influences $\vec{n}$ nor $\Delta E$. It only takes effect on $\Delta f$, which is equal to 1 in the main text where for a finite gap the Fermi energy lies between the two bands. Further, note that the evaluation of the $\delta(\omega-\frac{\Delta E}{\hbar})$ term results in the integral being reduced to a surface integration. Solving $\omega-\frac{\Delta E}{\hbar}=0$ for $k_{z}$, we find that due to the presence of the symmetry-enforced nodal plane $k_{z,0}=a(k_{x},k_{y})\omega+\mathcal{O}(\omega^{2})$ with some function $a(k_{x},k_{y})$, as long as the nodal plane is the only degeneracy of Eq. (1). Then similarly, we get the low frequency expansion $\partial_{z}\vec{n}\cdot\partial_{z}\vec{n}=b(k_{x},k_{y})\omega^{2}+\mathcal{O}(\omega^{3})$ by plugging in the previous $k_{z}$ expansion. With $d_{0}(\mathbf{k})=0$ (the case from the main text) and these considerations, Eq. (8) for $a=b=z$ becomes
@@
$\times\int\text{d}k_{x}\text{d}k_{y}h(k_{x},k_{y})\Delta f(k_{x},k_{y},k_{z}=k_{z,0}(\omega))+\mathcal{O}(\omega^{4})\;.$ (18)
@@
$\hskip 56.9055pt\times\left(63a^{2}+\frac{15a\left(f(\omega)-2a\varepsilon^{2}\right)^{2}}{\varepsilon^{2}f(\omega)}+\frac{35\left(f(\omega)-2a\varepsilon^{2}\right)^{4}}{36\varepsilon^{4}f(\omega)^{2}}\right)\;,$ (20)
@@
$\hskip 56.9055pt\times 2t_{\mathrm{c}}\cos(k_{z})-\cos(k_{z})+\mu$
@@
$\hskip 56.9055pt\times 2t_{0}\cos\left(\frac{k_{z}}{2}\right)$
@@
$\hskip 56.9055pt\times 2t_{0}\cos\left(\frac{k_{z}}{2}\right)$
@@
$\hskip 56.9055pt\times 4t_{\mathrm{c}}\sin\left(\frac{k_{x}}{2}\right)\sin(k_{z})\;,$
where $\bm{\sigma}$ is the vector of Pauli matrices. We added the $\cos(k_{z})$ term and the chemical potential $\mu$ in $d_{0}$ to move the Weyl points far below the Fermi energy such that their contribution to the optical conductivity is suppressed. For our calculation we use the parameters $t_{0}=0.4$ and $t_{\mathrm{c}}=0.1$ and we set the chemical potential to $\mu=-1.24$. The resulting band structure is shown in Fig. 3(a).
By expanding the Hamiltonian (21) around the high-symmetry point H $(\frac{4\pi}{3},0,\pi)^{\mathrm{T}}$ one obtains the low-energy model
@@
$+\left(\frac{\sqrt{3}}{2}t_{c}k_{z}^{3}-b\;3\sqrt{3}t_{c}k_{z}\right)\sigma_{z},$ (22)
which allows us to understand the dependence of the analytical solution on the parameters of the tight-binding model. Note, that for this we neglected further terms cubic in $k$, which are expected to only give small corrections in the prefactors. Furthermore, the energy offset caused by the dispersion and the chemical potential is neglected. This low-energy model corresponds to the a=1a=-1 phase discussed for the model in Eq. (1), which means that two additional Weyl points are present. In order to remove their contribution to the optical response, we change the phase to correspond to the $a=+1$ case by introducing the parameter $b$ and setting b=1b=-1. With this, the analytical behaviour shown in Fig. 3(b) is obtained.
@@
$\times\left.\left(\frac{2}{3}\sqrt{\frac{2}{3}}\left(f(\omega)-\frac{20}{f(\omega)}-4\right)^{3/2}+\frac{\left(f(\omega)-\frac{20}{f(\omega)}-4\right)^{5/2}}{6\sqrt{6}}+\sqrt{6}\sqrt{f(\omega)-\frac{20}{f(\omega)}-4}\right)\right]\;,$ (25)
@@
which shows that this low-frequency behaviour does not depend on the topological charge but is a characteristic feature of nodal planes. In particular, every Berry curvature component in this model vanishes, demonstrating again the important role of the quantum metric for the investigation of response properties of materials.
APPENDIX F Suitable space groups and materials
In order to study the low-frequency optical response of a nodal plane material experimentally, one can select systems with certain properties, which are expected to lead to a more reliable determination of the power-law behaviour. In general, there are two requirements that a nodal plane material must satisfy in order to show a clear $\omega^{3}$ power law in its optical conductivity:
The first requirement is that the Drude peak, originating from intraband transitions of the nodal-plane bands (and possibly other bands near the Fermi energy), should be narrow in frequency. In general, for a narrow Drude peak, clean systems are required, which should be measured at low temperature, while making sure that the crystal structure is stable at low temperatures and no magnetic order develops. Furthermore, for the Drude peak to be small, no other band features should be located at the Fermi level apart from the nodal plane and the density of states should be low. We would like to note, however, that in principle it is possible to subtract the Drude contribution from the optical conductivity, to reveal the interband response, as recently demonstrated for the multi-fold fermion compound RhSi [76].
The second requirement is a broad frequency range for which the cubic power-law behaviour can be measured. As for the first point, this is expected to be optimal in systems with no other bands (and most importantly topological crossings) at the Fermi level. In particular, the energy range in which the two bands of the nodal plane disperse away from it without intersecting with each other or other bands should be as large as possible. Since in the case of a topological nodal plane the two bands need to form Weyl points in order to compensate the topological charge, trivial nodal plane systems are expected to be better suited for this. All of the discussed requirements can ensure a large frequency range, in which the nodal plane signature is expected to be visible.
With these criteria, we find that relevant space groups (SGs) should have only one two-fold screw rotation (for a single nodal plane) and mirror and/or inversion symmetries (for a trivial charge). We list all SGs which can host single nodal planes in Tab. 1, where we also indicate, whether spin-orbit-coupling (SOC) can gap out the nodal plane and whether a topological charge is allowed for the specific SG. With this, one can perform a search for previously synthesized materials, which fit our criteria to a reasonable degree and which we propose to be suitable for the optical conductivity measurement. In particular, we suggest that Sb3Sr5 (SG 193) [73] and XXMo3S3 ($X$ = Rb, K, Cs) (SG 176) [25, 74] are the best candidates for this. The latter material group is particularly promising, since these materials have an almost flat trivial nodal plane as the only feature at the Fermi level. We further propose HfI3 (SG 193) [77] and K4P3 (SG 63) [78] as suitable materials.
SG NP position NP w. SOC top. NP
4 $P1\,2_{1}\,1$ $k_{y}=\pi$ allowed
11 $P1\,2_{1}/m\,1$ $k_{y}=\pi$ no
14 $P1\,2_{1}/c\,1$ $k_{y}=\pi$ no
17 $P222_{1}$ $k_{z}=\pi$ allowed
20 $C222_{1}$ $k_{z}=\pi$ allowed
26 $Pmc2_{1}$ $k_{z}=\pi$ no
29 $Pca2_{1}$ $k_{z}=\pi$ no
31 $Pmn2_{1}$ $k_{z}=\pi$ no
33 $Pna2_{1}$ $k_{z}=\pi$ no
36 $Cmc2_{1}$ $k_{z}=\pi$ no
51 $P2_{1}/m\,2/m\,2/a$ $k_{z}=\pi$ no
52 $P2/n\,2_{1}/n\,2/a$ $k_{z}=\pi$ no
53 $P2/m\,2/n\,2_{1}/a$ $k_{z}=\pi$ no
54 $P2_{1}/c\,2/c\,2/a$ $k_{z}=\pi$ no
63 $C2/m\,2/c\,2_{1}/m$ $k_{z}=\pi$ no
64 $C2/m\,2/c\,2_{1}/e$ $k_{z}=\pi$ no
76 $P4_{1}$ $k_{z}=\pi$ allowed
78 $P4_{3}$ $k_{z}=\pi$ allowed
91 $P4_{1}22$ $k_{z}=\pi$ allowed
95 $P4_{3}22$ $k_{z}=\pi$ allowed
169 $P6_{1}$ $k_{z}=\pi$ allowed
170 $P6_{5}$ $k_{z}=\pi$ allowed
173 $P6_{3}$ $k_{z}=\pi$ allowed
176 $P6_{3}/m$ $k_{z}=\pi$ no
178 $P6_{1}22$ $k_{z}=\pi$ allowed
179 $P6_{5}22$ $k_{z}=\pi$ allowed
182 $P6_{3}22$ $k_{z}=\pi$ allowed
185 $P6_{3}cm$ $k_{z}=\pi$ no
186 $P6_{3}mc$ $k_{z}=\pi$ no
193 $P6_{3}/m\,2/c\,2/m$ $k_{z}=\pi$ no
194 $P6_{3}/m\,2/m\,2/c$ $k_{z}=\pi$ no
Table 1: All space groups with single nodal planes. First and second column: Space group (SG) number and full symbol. Third column: Position of the nodal plane. Fourth column: Check mark denotes SGs, for which spin-orbit-coupling (SOC) splits the spin-degenerate nodal plane (without SOC) into two nodal planes, otherwise the nodal planes are gapped with SOC. Fifth column: In the absence of mirror symmetries, a nodal plane may carry a non-zero Chern number, denoted as “allowed” topological nodal plane.